### v1

# Propagation of Singularities for the Damped Stochastic Klein-Gordon Equation

Abstract
For the $1+1$ dimensional damped stochastic Klein-Gordon equation, we show that random singularities associated with the law of the iterated logarithm exist and propogate in the same way as the stochastic wave equation. This provides evidence for possible connections to microlocal analysis, ie. the exact regularity and singularities described in this paper should admit wavefront set type descriptions whose propagation is determined by the highest order terms of the linear operator. Despite the results being exactly the same as those of the wave equation, our proofs are significantly different than the proofs for the wave equation. Miraculously, proving our results for the critically damped equation implies them for the general equation, which significantly simplifies the problem. Even after this simplification, many important parts of the proof are significantly different than (and we think are more intuitive from the PDE viewpoint compared to) existing proofs for the wave equation.
1 Introduction
We study the regularity of the one-dimensional damped stochastic Klein-Gordon equation
$(\mathop{}\!\mathbin{\Box}+a\partial_{t}+m^{2})u(t,x)=\dot{W}(t,x),\quad t>0,x\in\mathbb{R},$ (1)
interpreted in the sense of [Wal86] subject to initial conditions $u(0,x)=\partial_{t}u(0,x)=0$. Here $a,m\in\mathbb{R}$ are known from the physics literature as the damping constant and mass, $\mathop{}\!\mathbin{\Box}=\partial_{tt}-\partial_{xx}$ is the d’Alembertian in one space dimension, and $\dot{W}$ is Gaussian space-time white noise defined on a complete probability space $(\Omega,\mathscr{F},\mathbb{P})$. Let $\{\mathscr{F}_{t}\}_{t\geq 0}$ denote the filtration generated by the noise $\dot{W}$. Our main results are the following Theorems.
Theorem 1.1 (Law of the iterated logarithm).
For any fixed $t_{0}\geq 0$ and $x_{0}\in\mathbb{R}$, there exists a constant $0<K_{1}<\infty$ depending on $(t_{0},x_{0})$ such that
$\limsup_{h\to 0^{+}}\frac{|u(t_{0}+h,x_{0}+h)-u(t_{0},x_{0})|}{\sqrt{h\log\log(1/h)}}=K_{1}\quad\text{a.s.}$ (2)
Theorem 1.2 (Modulus of continuity).
Fix $w_{0}\geq 0$ and $0<a_{1}<a_{2}$, and consider the line segment
$I=\left\{(t,x)\in\mathbb{R}_{+}\times\mathbb{R}:\tfrac{t-x}{\sqrt{2}}=w_{0},\tfrac{t+x}{\sqrt{2}}\in[a_{1},a_{2}]\right\}.$ (3)
Then, there exist constants $0<C_{1}<C_{2}<\infty$ depending on $(w_{0},a_{1},a_{2})$ such that for every sub-interval $J$ of $I$, there exists a constant $K_{J}\in[C_{1},C_{2}]$ such that
$\limsup_{h\to 0^{+}}\sup_{(t,x)\in J}\frac{|u(t+h,x+h)-u(t,x)|}{\sqrt{h\log(1/h)}}=K_{J}\quad\text{a.s.}$ (4)
Theorem 1.1 states that at a fixed space-time point, the increments (along a characteristic direction) are locally of order $\sqrt{h\log\log(1/h)}$, while Theorem 1.2 asserts that the uniform modulus of continuity for the increments are of a larger order, at a logarithmic level. Therefore, Theorems 1.1 and 1.2 together suggest that there exist random space-time points at which local increments are larger than those at a fixed point. Theorem 1.4 below justifies the existence of these random singularities and shows that they propagate along the other characteristic direction.
One of the ingredients for proving Theorem 1.4 is the following simultaneous upper bound for the law of the iterated logarithm. For notational convenience, we introduce a change of coordinates via a rotation by $45^{\circ}$:
$(w,z)=\left(\tfrac{t-x}{\sqrt{2}},\tfrac{t+x}{\sqrt{2}}\right),\qquad(t,x)=\left(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}\right).$ (5)
The $w$ and $z$ directions are the characteristic directions of the wave equation.
Theorem 1.3.
Fix $N>0$ and $0<a_{1}<a_{2}$. Then, there exists a constant $0<K_{2}<\infty$ such that for all fixed $z_{0}\in[a_{1},a_{2}]$,
$\mathbb{P}\left\{\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w+z_{0}}{\sqrt{2}}+h,\frac{-w+z_{0}}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w+z_{0}}{\sqrt{2}},\frac{-w+z_{0}}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}\leq K_{2}\text{ for all $w\in[0,N]$}\right\}=1.$ (6)
Theorem 1.4 (Existence and propagation of singularities).
Fix $w_{0}\geq 0$, $0<z_{0}<z_{0}^{\prime}$. Let $t_{0}=w_{0}/\sqrt{2}$. Then:
  1. (i)
    There exists an $\mathscr{F}_{t_{0}}$-measurable random variable $Z=Z(\omega)$ such that $z_{0}\leq Z\leq z_{0}^{\prime}$ a.s. and
    $\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w_{0}+Z}{\sqrt{2}}+h,\frac{-w_{0}+Z}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w_{0}+Z}{\sqrt{2}},\frac{-w_{0}+Z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}=+\infty\quad\text{a.s.}$ (7)
  2. (ii)
    If $Z$ is any such random variable, then
    $\mathbb{P}\left\{\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w+Z}{\sqrt{2}}+h,\frac{-w+Z}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w+Z}{\sqrt{2}},\frac{-w+Z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}=+\infty\text{ for all $w>w_{0}$}\right\}=1.$ (8)
Remark 1.5.
In general one can define $\mathop{}\!\mathbin{\Box}=b^{2}\partial_{tt}-c^{2}\partial_{xx}$, but a simple change of coordinates would allow one to produce the above results (with appropriately tilted characteristic lines) with the same calculations. We also note that all of our results can apply to the excited stochastic Klein-Gordon equation $(a<0)$, but with extra dependence on $T$ the time horizon in all of the constants in the estimates.
Remark 1.6.
Theorems 1.11.2 together imply that the $B^{s}_{\infty,\infty}$-Besov wavefront set [DRS22] of $u$ is empty if $s<\frac{1}{2}$ and is equal to $(\mathbb{R}_{+}\times\mathbb{R})\times(\mathbb{R}^{2}\setminus\{(0,0)\})$ if $s\geq\frac{1}{2}$. With some extra work, the same statements can be said about the more familiar $H^{s}$-Sobolev wavefront sets [Hin25, Definition 7.35, Proposition 7.36] of $u$ with the same regularity parameters. We will discuss the relevance of this in the literature review.
1.1 Literature review and motivation
The (Khinchin) law of the iterated logarithm (LIL, for short) was first stated by Khinchin in 1924 [Khi24] as a description of the fluctuations of a symmetric random walk. Since Brownian motion is the limit of scaled random walks, the LIL for Brownian motion [Khi33] can be interpreted as both regularity and large time fluctuation statements. We will concentrate on the regularity interpretation, for which we note that it has since been generalized to stochastic differential equations and Brownian sheets [Zim72, OP73]. Importantly, local versions of the LIL with a different power of $h$ in the denominator of (2) has been proven for fractional Brownian Motion and fractional Brownian sheets [Ore72, AX05], which can be seen as a reflection of their regularity being higher or lower than Brownian motion. All of these have also been generalized to Gaussian processes, Gaussian random fields, and stochastic partial differential equations (SPDEs) driven by appropriate objects [MR06, MWX13, LX22, HL25].
Another regularity property of Brownian motions, Levy’s modulus of continuity (MC) [Lév37], is apparently at odds with the LIL since it divides by a larger denominator compared to the statement of LIL. This has also been generalized to Gaussian processes and SPDEs [MR06, MWX13, LX19, LX23, HL25]. Notably, another way the LIL and MC are different is that the LIL fixes a point in parameter space while MC allows variation of the base point in parameter space. This indicates the existence of random points at which local increments are exceptionally large and fail to satisfy LIL. We will call such points LIL-singularities. Exceptional increments were first studied for the Brownian motion by Orey and Taylor [OT74] who showed that the set of LIL-singularities is dense, uncountable, and has Hausdorff dimension 1, and also studied different types of singularities. Similar results are known for Brownian sheets, general Gaussian processes, and SPDEs [Wal82, Wal86, KS00, KPX00, HK17]. In particular, Walsh [Wal82] studied the LIL-singularities of a 2-parameter Brownian sheet and proved that singularities along an axis direction propagate perpendicularly along the other axis direction. Walsh modified his argument to show that LIL-singularities of stochastic wave equation propagate along characteristic directions [Wal86]. These results have been extended to semi-fractional Brownian sheets [BM08] and more general stochastic wave equations [CN88, LX22].
Interpreting LIL singularities as local statements about behavior near a point hints at possible connections to microlocal analysis [Sat70, Hör71, Hör73, Vas11, Hin25], which encompasses the sharpest propagation of singularities literature for deterministic equations. In microlocal analysis, singularities are characterized using wavefront sets. Informally, a point $(x,\xi)\in\mathbb{R}^{n}\times(\mathbb{R}^{n}\setminus\{0\})$ (viewed as the cotangent bundle of $\mathbb{R}^{n}$) is an element of the (smooth) wavefront set of some Schwartz distribution $\phi$ if $\phi$ is not smooth in the direction $\xi$ at $x$. The celebrated paper [Hör78] proved that for solutions of evolution equations governed by some linear operator $P$, wavefront sets of initial conditions and forcing terms evolve according to the Hamiltonian flow of $P$, which is entirely determined by its highest order terms. The simplest application of this result is that equations of the form $(\mathop{}\!\mathbin{\Box}+P_{L})u=0$, where $P_{L}$ is an operator of order <2<2, propagate singularities in the same way as $\mathop{}\!\mathbin{\Box}u=0$. Notably, it also applies to operators with variable coefficients. With physical motivations, the definition of the wavefront set and the result of [Hör78] has been extended to Sobolev norms with arbitrary (including negative) regularity [SL82, Den82], and most recently, to Besov norms [DRS22]. For concrete examples, the $H^{s}$-Sobolev wavefront set (here $H^{s},s\in\mathbb{R}$ is the standard $L^{2}$-based Sobolev space) of $\delta_{x}$ is empty if $s<-\frac{n}{2}$ and is equal to the nonzero cotangent space at $x$ if $s\geq-\frac{n}{2}$, and the $H^{s}$-Sobolev wavefront set of $\mathbb{1}_{\mathcal{D}}$ for a smooth domain $\mathcal{D}\subset\mathbb{R}^{n}$ is empty for $s<\frac{1}{2}$ and is equal to the nonzero normal bundle of $\partial_{D}$ if $s\geq\frac{1}{2}$ [Hin25, Example 7.37]. Due to the diffeomorphism invariance of wavefront sets, all of this theory can be extended to smooth manifolds, which has been used in mathematical general relativity for both stability and singularity generation [HV18, Hin21, HHV21].
We conjecture that as regularity statements, an order $\beta$ modulus of continuity($h^{\beta}\log\frac{1}{h}$ in the denominator of Theorem 1.2) or law of iterated logarithm($h^{\beta}\log\log\frac{1}{h}$ in the denominator of Theorem 1.1) holding at some point $x$ for some distribution $\phi$ can be characterized as the cotangent space at $x$ not belonging to a $\beta$-MC wavefront set” or $\beta$-LIL wavefront set”, while $\beta$-LIL singularities would correspond to the existence of a random element of $\mathbb{R}^{2n}$, which we can interpret as a random $\beta$-LIL wavefront set”. The basis for our conjecture is the well-known correspondence between small scale and high frequency behavior, which is the intuition behind the Sobolev and Besov wavefront set definitions. Furthermore, we also believe that once the definition of a $\beta$-MC/LIL wavefront set has been written down, it would be valid for $\beta$ negative, which would allow for investigation of these properties beyond random fields. We also conjecture that [Hör78] can be extended to these MC/LIL-wavefront sets(the LIL extension would be more interesting considering Theorem 1.4). Affirmative answers to our conjectures would not only greatly extend all current results about MC, LIL, and LIL-type singularities, but also can be relevant for many other areas of continuous probability. For instance, it would allow for more precise regularity descriptions and even speak of pointwise properties of many objects of active study such as white noise, Gaussian free fields, LQG measures, and solutions of renormalized SPDEs, all of which are known to have negative regularity.
[…]
$(\partial_{t}-\mathop{}\!\mathbin{\bigtriangleup}-m^{2})v(t,x)=\dot{W}(t,x)$ (9)
will not propagate its singularities, which are known to exist in $1+1$ dimensions [HL25]. Our main results provide heuristics for this in $1+1$ dimensions using the Smolchowski- Kramers approximation
$(\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx}-m^{2})v_{\mu}(t,x)=\dot{W}(t,x),\quad\mu\in(0,1]$ (10)
where one will see (9) as $\mu\downarrow 0$ (we refer interested readers to [CF06, CS16, CX23] for the stability of this approximation and related topics). For $\mu>0$, one can easily adopt the contents of this paper to see that Theorems 1.11.4 all apply to $v_{\mu}$ with appropriately tilted characteristic lines. The effect of tilting the characteristic lines as $\mu$ gets smaller is that for any singularity of $v_{\mu}$ at some point $(t_{0},x_{0})$, the propagated singularity of part (ii) of Theorem 1.4 is at $(t_{0}+\varepsilon,x_{0}+\frac{\varepsilon}{\mu})$ at time $t_{0}+\varepsilon$, which notably goes to infinity as $\mu\downarrow 0$. This is consistent with the PDE heuristic that hyperbolic operators have finite speed of propagation while parabolic operators have infinite speed of propagation, so $\mathop{}\!\mathbin{\Box}_{\mu}:=\mu^{2}\partial_{tt}-\partial_{xx}$ indexes a family of hyperbolic operators whose speed of propagation goes to infinity as $\mu\downarrow 0$.
1.2 Paper organization and notation
The rest of the paper will be organized as follows. In section 2 we state our Fourier transform convention and prove pointwise and increment second moment bounds for our solution $u$. In section 3, we show that proving any of our main theorems in the case of the critically damped equation actually implies the same theorem for all damped equations. The rest of the paper is then dedicated to proving the main theorems for the critically damped equations. Notably, the proofs of Theorem 1.3 and the propagation part of Theorem 1.4 are significantly different compared to existing proofs of analogous results for the wave equation, despite the simplification of section 3. In fact, one can see directly in the propagation proof that damping does not erase the singularity.
We list some notation we will employ for the rest of the paper.
  • Sets: $\mathbb{R}_{+}=(0,\infty)$, $\mathbb{N}=\{0,1,2,\dots\}$, $\mathbb{N}_{+}=\{1,2,\dots\}$
  • $a\wedge b=\min\{a,b\}$; $\mathbb{1}_{A}$ denotes indicator function of a set AA
  • $f(x)\propto g(x)$ means that there exists $0<C<\infty$ such that $f(x)=Cg(x)$ for all $x$, and $f(x)\lesssim g(x)$ means that there exists $C<\infty$ such that $f(x)\leq Cg(x)$ for all $x$. In one calculation, we may write these in succession without mentioning that the constants change from line to line, and we will only note dependence of the constants on certain parameters if we deem it relevant to do so.
2 Preliminaries
For notational clarity, we choose a Fourier transform convention and state some basic theorems in it.
Definition 2.1.
For $f\in L^{2}(\mathbb{R})$, define its Fourier transform by
$\mathcal{F}[f](\xi)=\hat{f}(\xi):=\int_{\mathbb{R}}f(x)e^{-ix\xi}dx.$
The following are standard results.
  1. 1.
    Translation: $\mathcal{F}[f(\cdot-a)](\xi)=e^{-ia\xi}\hat{f}(\xi)$.
  2. 2.
[…]
  3. 3.
    Parseval-Plancherel identity: $\braket{f,g}_{L^{2}(\mathbb{R})}=\frac{1}{2\pi}\braket{\hat{f},\hat{g}}_{L^{2}(\mathbb{R})}$.
  4. 4.
    Convolution: define $f*g(x):=\int_{\mathbb{R}}f(x-y)g(y)dy$. Then $\mathcal{F}[f*g](\xi)=\hat{f}(\xi)\hat{g}(\xi).$
The solution to (1) is
$u(t,x)=\int_{0}^{t}\int_{-\infty}^{\infty}G(t-s,x-y)W(ds,dy)$ (11)
where $G$ denotes the fundamental solution to $\mathop{}\!\mathbin{\Box}+a\partial_{t}+m^{2}$ and the integral on the right hand side is understood in the sense of Walsh [Wal86]. The existence and regularity of $G$ is an elementary PDE exercise. Taking the Fourier transform in space, one finds that $\hat{G}(t,\xi)$, the Fourier transform of $G(t,\cdot)$, satisfies
[…]
Observe that the small frequency regime exists if and only if $\frac{a^{2}}{4}-m^{2}>0$. Note that $G$ is non-elementary except in the critial damping case $m^{2}=a^{2}/4$ (see (16) below). The following estimates will be useful.
Lemma 2.2.
Let $u$ be defined as in (11). Fix $0<T<\infty$. Then, for all $0<s<t\leq T$ and $x\in\mathbb{R}$,
$\mathbb{E}[u(t,x)^{2}]\lesssim t^{2}$
where the implicit constant depends only on $T$, $a$, $m$.
Proof.
Using Itô–Walsh isometry, the change of variables $s\mapsto t-s$, $y\mapsto x-y$, and Plancherel’s theorem, we can write
$\mathbb{E}[u(t,x)^{2}]=\int_{0}^{t}\int_{-\infty}^{\infty}|G(t-s,x-y)|^{2}dyds=\int_{0}^{t}\int_{-\infty}^{\infty}|G(s,y)|^{2}dyds\propto\int_{0}^{t}\int_{-\infty}^{\infty}|\hat{G}(s,\xi)|^{2}d\xi ds.$
Then, by (12),
$\mathbb{E}[u(t,x)^{2}]\propto$ $\int_{0}^{t}\int_{-\infty}^{\infty}\left[e^{-as}\frac{\sin^{2}\left(s\sqrt{\xi^{2}+m^{2}-\frac{a^{2}}{4}}\right)}{\xi^{2}+m^{2}-\frac{a^{2}}{4}}\mathbb{1}_{\set{\xi^{2}>\frac{a^{2}}{4}-m^{2}}}\right$.
$\qquad\left.+e^{-as}\frac{\sinh^{2}{\left(t\sqrt{m^{2}-\frac{a^{2}}{4}-\xi^{2}}\right)}}{m^{2}-\frac{a^{2}}{4}-\xi^{2}}\mathbb{1}_{\set{\xi^{2}<\frac{a^{2}}{4}-m^{2}}}\right]d\xi ds$
$=$ $\,I_{1}+I_{2},$
where $K^{2}=(m^{2}-\frac{a^{2}}{4})\wedge 0$,
$I_{1}:=\int_{0}^{t}\int_{|\xi|<K}e^{-as}\frac{\sinh^{2}{\left(s\sqrt{K^{2}-\xi^{2}}\right)}}{K^{2}-\xi^{2}}d\xi ds,\quad I_{2}:=\int_{0}^{t}\int_{|\xi|>K}e^{-as}\frac{\sin^{2}(s\sqrt{\xi^{2}-K^{2}})}{\xi^{2}-K^{2}}.$
For $I_{1}$, we have
[…]
The first $\lesssim$ above used the fact that $\frac{\sinh(x)}{x}$ is uniformly bounded over any compact interval $I\subset\mathbb{R}$.
For $I_{2}$, we choose and fix some constant $L>K$, say $L=K+1$, and further decompose $I_{2}=J_{1}+J_{2},$ where
[…]
Lemma 2.3.
For $0\leq s<t\leq T$ and $x\in\mathbb{R}$, we have
$\mathbb{E}[|u(t,x)-u(s,x)|^{2}]\lesssim s(t-s)$
where the implicit constant depends only on $T$, $a$, $m$.
Proof.
We have
$u(t,x)-u(s,x)=\int_{0}^{s}\int_{\mathbb{R}}G(t-\tau,x-y)-G(s-\tau,x-y)W(d\tau,dy)+\int_{s}^{t}\int_{\mathbb{R}}G(t-\tau,x-y)W(d\tau,dy),$
giving us
[…]
$+\int_{s}^{t}\int_{\mathbb{R}}|\hat{G}(t-\tau,\xi)|^{2}d\xi d\tau=:I_{1}+I_{2}.$
By a change of variable, the moment bound in Lemma 2.2 gives $I_{2}\lesssim(t-s)^{2}\wedge 1$.
For $I_{1}$, again we use $K^{2}=(m^{2}-\frac{a^{2}}{4})\wedge 0$, choose and fix a constant $L>K$, say $L=K+1$, and apply the decomposition
$I_{1}=\int_{0}^{s}\int_{\mathbb{R}}(\cdots)=\int_{0}^{s}\int_{|\xi|<K}(\cdots)+\int_{0}^{s}\int_{K<|\xi|<L}(\cdots)+\int_{0}^{s}\int_{|\xi|>L}(\cdots)=:J_{1}+J_{2}+J_{3}.$
For $J_{1}$, we use the following application of the triangle inequality
$|e^{-x}f(t)-e^{-y}f(s)|\leq|e^{-x}(f(t)-f(s))|+|(e^{-x}-e^{-y})f(s)|,$
together with the elementary inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$, to deduce the following:
$J_{1}\propto$ $\int_{0}^{s}\int_{|\xi|<K}\left[e^{-\frac{a(t-\tau)}{2}}\frac{\sinh\left((t-\tau)\sqrt{K^{2}-\xi^{2}}\right)}{\sqrt{K^{2}-\xi^{2}}}-e^{-\frac{a(s-\tau)}{2}}\frac{\sinh\left((s-\tau)\sqrt{K^{2}-\xi^{2}}\right)}{\sqrt{K^{2}-\xi^{2}}}\right]^{2}d\xi d\tau$
[…]
The third inequality used the mean value theorem and the fact that both $\exp$ and $\cosh$ are increasing on $\mathbb{R}_{+}$. The singularities of the integrand at $\xi=0,\pm K$ are integrable by an elementary calculus argument.
Similarly, we have
$J_{2}\propto$ $\int_{0}^{s}\int_{K<|\xi|<L}\left[e^{-\frac{a(t-\tau)}{2}}\frac{\sin\left((t-\tau)\sqrt{\xi^{2}-K^{2}}\right)}{\sqrt{K^{2}-\xi^{2}}}-e^{-\frac{a(s-\tau)}{2}}\frac{\sin\left((s-\tau)\sqrt{K^{2}-\xi^{2}}\right)}{\sqrt{\xi^{2}-K^{2}}}\right]^{2}d\xi d\tau$
[…]
$\lesssim$ $\int_{0}^{s}\int_{K<|\xi|<L}\Bigg{[}\sqrt{\xi^{2}-K^{2}}(t-s)+\sin((t-\tau)\sqrt{\xi^{2}-K^{2}})(t-s)\frac{a}{2}\Bigg{]}^{2}\frac{d\xi d\tau}{\xi^{2}-K^{2}}$
[…]
Again the singularity at $K$ is integrable by elementary arguments.
For $J_{3}$, we split it into two parts as we did above. For the first term, use the trigonometric identity $\sin(A)-\sin(B)=2\cos(\frac{A+B}{2})\sin(\frac{A-B}{2})$ and bound the $\cos$ factor by 1. For the second term, bound $\sin$ by 1 and apply mean value theorem to the difference of the exponential functions. This yields the following:
[…]
$\leq$  $\,2\int_{0}^{s}\int_{|\xi|>L}\Bigg{[}4\sin^{2}((t-s)\sqrt{\xi^{2}-K^{2}})+a^{2}(t-s)^{2}\sin^{2}((t-\tau)\sqrt{\xi^{2}-K^{2}})\Bigg{]}\frac{d\xi d\tau}{\xi^{2}-K^{2}}$
$\lesssim$ $\int_{0}^{s}\int_{|\xi|>L}\frac{\sin^{2}((t-s)\sqrt{\xi^{2}-K^{2}})}{\xi^{2}-K^{2}}d\xi d\tau+(t-s)^{2}\int_{0}^{s}\int_{|\xi|>L}\frac{\sin^{2}((t-\tau)\sqrt{\xi^{2}-K^{2}})}{\xi^{2}-K^{2}}d\xi d\tau=:S_{1}+S_{2}.$
Since $L>K$, we have $S_{2}\lesssim s(t-s)^{2}$. For $S_{1}$, we use the change of variable $\xi\mapsto\xi/(t-s)$, giving us
$S_{1}=$ $\,s(t-s)\int_{(t-s)L}^{\infty}\frac{\sin^{2}(\frac{1}{2}\sqrt{\xi^{2}-(t-s)^{2}K^{2}})}{\xi^{2}-(t-s)^{2}K^{2}}d\xi$
$\leq$ $\,s(t-s)\int_{(t-s)L}^{\infty}\frac{1\wedge(\xi^{2}-(t-s)^{2}K^{2})}{\xi^{2}-(t-s)^{2}K^{2}}d\xi$
$\leq$ $\,s(t-s)\left(\int_{(t-s)L}^{2TL}1d\xi+\int_{2TL}^{\infty}\frac{1}{\xi^{2}/2}d\xi\right)\propto s(t-s).$
This finishes the proof of Lemma 2.3. ∎
Lemma 2.4.
For any $t>0$ and $x,y\in\mathbb{R}$, we have
$\mathbb{E}[|u(t,x)-u(t,y)|^{2}]\lesssim t|x-y|,$
where the implicit constant depends only on $T$, $a$, $m$.
Proof.
We can write
[…]
$\leq$ $\int_{0}^{t}\int_{-\infty}^{\infty}\left(4\wedge(|x-y|^{2}|\xi|^{2})\right)\hat{G}(s,\xi)^{2}d\xi ds$
$\lesssim$ $\int_{0}^{t}\int_{0}^{1/|x-y|}|x-y|^{2}|\xi|^{2}\hat{G}(s,\xi)^{2}d\xi ds+\int_{0}^{t}\int_{1/|x-y|}^{\infty}|\hat{G}(s,\xi)|^{2}d\xi ds.$
The bound $|\hat{G}(s,\xi)|^{2}\lesssim\frac{1}{\xi^{2}}$ for all $\xi>0$ gives
$\int_{0}^{t}\int_{1/|x-y|}^{\infty}|\hat{G}(s,\xi)|^{2}d\xi ds\lesssim\int_{0}^{t}\int_{1/|x-y|}^{\infty}\frac{1}{\xi^{2}}d\xi ds\leq t|x-y|.$
If |xy|>1|x-y|>1, then $|\xi|<1/|x-y|$ implies $|\xi|<1$. This together with $|\hat{G}(s,\xi)|\lesssim 1$ for $|\xi|<1$ gives
$\int_{0}^{t}\int_{0}^{1/|x-y|}|x-y|^{2}|\xi|^{2}\hat{G}(s,\xi)^{2}d\xi ds\lesssim|x-y|^{2}\int_{0}^{t}\int_{0}^{1/|x-y|}d\xi ds=t|x-y|.$
Finally, if $|x-y|\leq 1$, we use $\int_{0}^{1/|x-y|}=\int_{0}^{1}+\int_{1}^{1/|x-y|}.$ The first estimate and $|\xi|<1$ then gives
$\int_{0}^{t}\int_{0}^{1}|x-y|^{2}|\xi|^{2}\hat{G}(s,\xi)^{2}d\xi ds\lesssim|x-y|^{2}t^{2}.$
For $|\xi|>1$, we again use $|\hat{G}(s,\xi)|^{2}\lesssim\frac{1}{\xi^{2}}$, so
$\int_{0}^{t}\int_{1}^{1/|x-y|}|x-y|^{2}|\xi|^{2}\hat{G}(s,\xi)^{2}d\xi ds\lesssim t(|x-y|^{2}-|x-y|)\lesssim t|x-y|.$
This finishes the proof of Lemma 2.4. ∎
Combining Lemma 2.3 and Lemma 2.4, and applying the Kolmogorov continuity theorem, we obtain:
Proposition 2.5.
For any $T>0$, there exists $C>0$ such that
$\mathbb{E}[|u(t,x)-u(s,y)|^{2}]\leq C(|t-s|+|x-y|)$ (13)
uniformly for all $s,t\in[0,T]$ and $x,y\in\mathbb{R}$. Hence, the sample function $(t,x)\mapsto u(t,x)$ is a.s. continuous on $0,\infty)\times\mathbb{R}$ and locally Hölder of order $1/2$- in both $t$ and $x$.
3 Reduction to the critical damping case
We now explain how the proofs of the main theorems can be reduced to the critical damping case $m^{2}=a^{2}/4$. Note that this simplification immediately implies all of the main theorems if $a=0$ by [Wal86, CN88, LX22], so for all future sections we will assume $a>0$ to avoid triviality. First, re-write equation (1) as:
$\begin{cases}(\mathop{}\!\mathbin{\Box}+a\partial_{t}+a^{2}/4)u(t,x)=\dot{W}(t,x)+b(u(t,x)),\quad t>0,x\in\mathbb{R},\\ u(0,x)=\partial_{t}u(0,x)=0,\end{cases}$ (14)
where $b:\mathbb{R}\to\mathbb{R}$ is the function defined by
$b(x)=(a^{2}/4-m^{2})x.$
Since $b$ is globally Lipschitz and grows linearly, standard existence and uniqueness theory implies that the solution $u$ to (1) is also the unique mild solution to (14) which solves the integral equation
$u(t,x)=\int_{0}^{t}\int_{\mathbb{R}}\Gamma(t-s,x-y)W(ds,dy)+\int_{0}^{t}\int_{\mathbb{R}}\Gamma(t-s,x-y)b(u(s,y))dyds,\quad t>0,x\in\mathbb{R},$ (15)
where $\Gamma$ is the fundamental solution to $\mathop{}\!\mathbin{\Box}+a\partial_{t}+a^{2}/4$. Notice that in this case, we have by (12)
$\hat{\Gamma}(t,\xi)=\frac{e^{-at/2}\sin(t|\xi|)}{|\xi|}\quad\text{and hence}\quad\Gamma(t,x)=\frac{e^{-at/2}}{2}\mathbb{1}_{\{|x|<t\}}.$ (16)
As mentioned in the preliminaries, this special case is referred to as critical damping and significantly simplifies our problem. To show that the simplification is valid, we write (15) as
$u(t,x)=u_{C}(t,x)+u_{L}(t,x),\text{ where }$ (17)
$u_{C}(t,x)=\int_{0}^{t}\int_{\mathbb{R}}\Gamma(t-s,x-y)W(ds,dy),$ (18)
$u_{L}(t,x)=(a^{2}/4-m^{2})\int_{0}^{t}\int_{\mathbb{R}}\Gamma(t-s,x-y)u(s,y)dyds.$ (19)
In particular, $u_{C}$ is the mild solution to (1) with $m^{2}=a^{2}/4$. In the following lemma, we show that $u_{L}$ is a.s. locally Lipschitz (along characteristic direction).
Lemma 3.1.
For any fixed compact set $F\subset[0,\infty)\times\mathbb{R}$, there exists $C>0$ such that
$\sup_{h\in[0,1]}\sup_{(t,x)\in F}\frac{|u_{L}(t+h,x+h)-u_{L}(t,x)|}{h}\leq C\sup_{(s,y)\in\mathscr{C}(F)}|u(s,y)|<+\infty\quad\text{a.s.}$ (20)
where $\mathscr{C}(F)=\bigcup_{h\in[0,1]}\bigcup_{(t,x)\in F}\{(s,y)\in[0,\infty)\times\mathbb{R}:|x+h-y|\leq t+h-s\}$.
Proof.
We first apply absolute value to the integrand to see that
$|u_{L}(t+h,x+h)-u_{L}(t,x)|$
$\leq|\tfrac{a^{2}}{4}-m^{2}|\int_{\mathbb{R}_{+}}\int_{\mathbb{R}}\left|\Gamma(t+h-s,x+h-y)\mathbb{1}_{\{|x+h-y|\leq t+h-s\}}-\Gamma(t-s,x-y)\mathbb{1}_{\{|x-y|\leq t-s\}}\right||u(s,y)|dyds$
$\leq|\tfrac{a^{2}}{4}-m^{2}|\sup_{(s,y)\in\mathscr{C}(F)}|u(s,y)|\,(I_{1}+I_{2}+I_{3}),$
where
$I_{1}=\int_{0}^{t}\int_{x-(t-s)}^{x+(t-s)}|\Gamma(t+h-s,x+h-y)-\Gamma(t-s,x-y)|dyds,$
$I_{2}=\int_{0}^{t}\int_{x+(t-s)}^{x+h+(t+h-s)}\Gamma(t+h-s,x+h-y)dyds,$
$I_{3}=\int_{t}^{t+h}\int_{x+h-(t+h-s)}^{x+h+(t+h-s)}\Gamma(t+h-s,x+h-y)dyds.$
Then, we may use (16) to estimate these three terms easily:
$I_{1}$ $=\frac{1}{2}\int_{0}^{t}\int_{x-(t-s)}^{x+(t-s)}(e^{-a(t-s)/2}-e^{-a(t+h-s)})dyds=(1-e^{-ah/2})\int_{0}^{t}(t-s)e^{-a(t-s)/2}ds\lesssim h,$
$I_{2}$ $=\frac{1}{2}\int_{0}^{t}\int_{x+(t-s)}^{x+(t-s)+2h}e^{-a(t+h-s)/2}dyds\leq h\int_{0}^{t}e^{-a(t-s)/2}ds\leq h,$
$I_{3}=\frac{1}{2}\int_{t}^{t+h}\int_{x-(t-s)}^{x+(t-s)+2h}e^{-a(t+h-s)}dyds\leq h^{2}.$
Finally, we finish the proof using the property that $u$ is a.s. bounded on the compact set $\mathscr{C}(F)$, due to the a.s. continuity of $u$ (see Proposition 2.5). ∎
3.1 Reduction of Theorem 1.1
Fix $t_{0}\geq 0$ and $x_{0}\in\mathbb{R}$. Assume that Theorem 1.1 holds for $u_{C}$, i.e., there exists a constant $0<K_{1}<\infty$ such that $\mathbb{P}(\Omega_{1})=1$, where
$\Omega_{1}=\left\{\limsup_{h\to 0^{+}}\frac{|u_{C}(t_{0}+h,x_{0}+h)-u_{C}(t_{0},x_{0})|}{\sqrt{h\log\log(1/h)}}=K_{1}\right\}.$
Thanks to Lemma 3.1, we may take $F=\{(t_{0},x_{0})\}$ and find an event $\Omega_{2}$ with $\mathbb{P}(\Omega_{2})=1$ on which (20) holds. This implies that for all $\omega\in\Omega_{2}$,
$\limsup_{h\to 0^{+}}\frac{|u_{L}(t_{0},x_{0}+h)-u_{L}(t_{0},x_{0})|}{\sqrt{h\log\log(1/h)}}=0.$
Hence, we may use (17) and the elementary inequality
$\limsup|f(x)|-\limsup|g(x)|\leq\limsup|f(x)+g(x)|\leq\limsup|f(x)|+\limsup|g(x)|$ (21)
(provided $\limsup|g(x)|<+\infty$) to deduce that (2) holds on the event $\Omega_{1}\cap\Omega_{2}$ which has probability 1.
3.2 Reduction of Theorem 1.2
Fix $w_{0}\geq 0$ and $0<a_{1}<a_{2}$. Assume that Theorem 1.2 holds for $u_{C}$, i.e., we can find constants $0<C_{1}<C_{2}<\infty$ such that for every sub-interval $J$ of $I$, there exists a constant $K_{J}\in[C_{1},C_{2}]$ such that $\mathbb{P}(\Omega_{J})=1$, where
$\Omega_{J}=\left\{\limsup_{h\to 0^{+}}\sup_{(t,x)\in J}\frac{|u(t+h,x+h)-u(t,x)|}{\sqrt{h\log(1/h)}}=K_{J}\right\}.$
Taking $F=I$ in Lemma 3.1, we can find an event $\Omega_{3}$ with $\mathbb{P}(\Omega_{3})=1$ on which (20) holds. It follows that for all $\omega\in\Omega_{3}$,
$\limsup_{h\to 0^{+}}\sup_{(t,x)\in I}\frac{|u_{L}(t+h,x+h)-u_{L}(t,x)|}{\sqrt{h\log(1/h)}}=0.$
Then, apply (17) and (21) to see that (4) holds on the event $\Omega_{J}\cap\Omega_{3}$ which has probability 1.
3.3 Reduction of Theorem 1.3
In order to simplify notations, write
$\ell(w,z)$ $=\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w+z}{\sqrt{2}}+h,\frac{-w+z}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w+z}{\sqrt{2}},\frac{-w+z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}},$
$\ell_{i}(w,z)$ $=\limsup_{h\to 0^{+}}\frac{|u_{i}\big{(}\frac{w+z}{\sqrt{2}}+h,\frac{-w+z}{\sqrt{2}}+h\big{)}-u_{i}\big{(}\frac{w+z}{\sqrt{2}},\frac{-w+z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}},\quad i=C,L.$
Fix $z_{0}>0$ and assume that Theorem 1.3 holds for $u_{C}$. This implies that there exists a constant $0<K_{2}<\infty$ such that $\mathbb{P}(A_{N})=1$ for all $N\in\mathbb{N}_{+}$, where
$A_{N}=\left\{\ell_{C}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\text{ for all $w\in[0,N]$}\right\}.$
An application of Lemma 3.1 with
$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[0,N],z=z_{0}\right\}$
shows that there is an event $B_{N}$ with $\mathbb{P}(B_{N})=1$ on which (20) holds. It is clear that
$B_{N}\subset\left\{\ell_{L}(w,z_{0})=0\text{ for all $w\in[0,N]$}\right\}.$
It follows from (17) and (21) that on the event $A_{N}\cap B_{N}$, which has probability 1, we have
$\ell(w,z_{0})\leq\ell_{C}(w,z_{0})+\ell_{L}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\quad\text{for all $w\in[0,N]$.}$
Finally, we may take $\bigcap_{N\in\mathbb{N}_{+}}(A_{N}\cap B_{N})$ as the desired event for which (6) holds.
3.4 Reduction of Theorem 1.4
Fix $w_{0}\geq 0$, $0<z_{0}<z_{0}^{\prime}$ and let $t_{0}=w_{0}/\sqrt{2}$. Assume that part (i) Theorem 1.4 holds for $u_{C}$. Then we can find an $\mathscr{F}_{t_{0}}$-measurable random variable $Z=Z(\omega)$ such that $\mathbb{P}(\Omega_{4})=1$, where $\Omega_{4}$ denotes the event
$\Omega_{4}=\{Z\in[z_{0},z_{0}^{\prime}]\text{ and }\ell_{C}(w_{0},Z)=+\infty\}.$
Taking
$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w=w_{0},z\in[z_{0},z_{0}^{\prime}]\right\}$
in Lemma 3.1, we can find another event $\Omega_{5}$ with $\mathbb{P}(\Omega_{5})=1$ on which (20) holds. In particular, this implies that
$\Omega_{5}\subset\{\ell_{L}(w_{0},z)=0\text{ for all $z\in[z_{0},z_{0}^{\prime}]$}\}.$
Therefore, $\Omega_{4}\cap\Omega_{5}$ is an event with probability 1 on which
$\ell(w,Z)\geq\ell_{C}(w,Z)-\ell_{L}(w,Z)=+\infty,$
thanks to (17) and (21). This proves that part (i) of Theorem 1.4 holds for $u$.
Let $Z$ be any $\mathscr{F}_{t_{0}}$-measurable random variable with $\mathbb{P}(\Omega_{6})=1$, where
$\Omega_{6}=\{Z\in[z_{0},z_{0}^{\prime}]\text{ and }\ell_{C}(w_{0},Z)=+\infty\},$
and assume that part (ii) of Theorem 1.4 holds for $u_{C}$. In particular, this assumption implies that $\mathbb{P}(E_{N})=1$ for all $N>w_{0}$, where
$E_{N}=\{\ell_{C}(w,Z)=+\infty\text{ for all $w\in[w_{0},N]$}\}.$
For each $N>w_{0}$, we may apply Lemma 3.1 with
$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[w_{0},N],z\in[z_{0},z_{0}^{\prime}]\right\}$
to see there is an event $G_{N}$ with $\mathbb{P}(G_{N})=1$ on which (20) holds. In particular,
$G_{N}\subset\{\ell_{L}(w,Z)=0\text{ for all $w\in[w_{0},N]$}\}.$
It follows that on $E_{N}\cap G_{N}$,
$\ell(w,Z)\geq\ell_{C}(w,Z)-\ell_{L}(w,Z)=+\infty\quad\text{for all $w\in[w_{0},N]$.}$
Hence, $\bigcap_{N\in\mathbb{N},N>w_{0}}(E_{N}\cap G_{N})$ is an event with probability 1 on which (8) holds. Therefore, part (ii) of Theorem 1.4 holds for $u$.
4 Proof of Theorems 1.1 and 1.2
Denote by $\tilde{F}[f](\tau,\xi)$ the space-time fourier transform of an arbitrary $f(t,x)$ and $\tilde{\Gamma}(\tau,\xi):=\tilde{F}[\Gamma(\cdot,\circ)](\tau,\xi)$.
Lemma 4.1.
Fix $T>0$. Define the random field $v(t,x)$ by
[…]
and the constrained random field $v(A,t,x)$ for a Borel set $A\subset[0,\infty)$ by
[…]
Then, $v(A,t,x)$ is a centered Gaussian random field such that $v(A,\cdot)$ and $v(B,\cdot)$ are independent whenever $A$ and $B$ are disjoint. Moreover, there exists $a_{0}>0$ such that for all $a_{0}\leq a<b\leq\infty$, $s,t\in[0,T]$ and $x,y\in\mathbb{R}$, we have
$\mathbb{E}[\{v([a,b),t,x)-v([a,b),s,y)-v(t,x)+v(s,y)\}^{2}]^{1/2}$
$\lesssim a^{2}(|t-s|+|x-y|)+b^{-2}$
and
$\mathbb{E}\left[(v([0,a_{0}],t,x)-v([0,a_{0}],s,y)\right]^{1/2}\lesssim|t-s|+|x-y|.$
We will need some elementary lemmas for the proof of Lemma 4.1.
Lemma 4.2.
Define $F(t,x,\tau,\xi):=\tilde{F}[\Gamma(t-\cdot,x-\circ)\mathbb{1}_{[0,t]}(\cdot)](\tau,\xi)$. For any $x,x^{\prime}\in\mathbb{R}$, $0<t,t^{\prime}<T$, the following holds.
  1. (a)
    $|F(t,x,\tau,\xi)|\lesssim\frac{1}{1+|\xi|}\left(\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}\right)$.
  2. (b)
    Given another $0<t^{\prime}<T$ and $x^{\prime}\in\mathbb{R}$,
    $|F(t,x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|\lesssim(|t-t^{\prime}|+|x-x^{\prime}|)\left[\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}\right].$
Proof.
This lemma is the analogue of [DMX15, Lemma 9.4] (with the results listed in reverse for convenience), and the proof will go similarly. By an elementary calculation, we have
$F(t,x,\tau,\xi)=$ $\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[f(t,z_{+})-f(t,z_{-})],\text{ where }$
$f(t,z):=$ $\frac{e^{tz}-1}{z},\quad z_{\pm}:=i\tau-\frac{a}{2}\pm i\xi.$ (24)
For the proof of part (a), (4) gives
$|F(t,x,\tau,\xi)|=\frac{|\phi_{t}(y_{+})-\phi_{t}(y_{-})|}{2|y_{+}-y_{-}|},\text{ where }\phi_{t}(y)=\frac{e^{it(y+i\frac{a}{2})-1}}{i(y+i\frac{a}{2})}.$
Observe that since $y_{+}-y_{-}=(y_{+}+i\frac{a}{2})-(y_{-}+i\frac{a}{2})$, we have by the substitution $\tilde{y}_{\pm}=y_{\pm}+i\frac{a}{2}$
$|F(t,x,\tau,\xi)|=\frac{|\tilde{\phi}_{t}(\tilde{y}_{+})-\tilde{\phi}_{t}(\tilde{y}_{-})|}{|\tilde{y}_{+}-\tilde{y}_{-}|},\text{ where }\tilde{\phi}_{t}(\tilde{y})=\frac{e^{it\tilde{y}}-1}{\tilde{y}}$.
Note here that $\tilde{\phi}_{t}$ is a map $\mathbb{C}\supset\mathbb{R}+i\frac{a}{2}\to\mathbb{C}$, so we need not worry about $\tilde{y}=0$. However, it is easy to show using the elementary arguments that
$\sup_{(t,\tilde{y})\in[0,T]\times\mathbb{R}+i\frac{a}{2}}\max(\tilde{\phi}_{t}(\tilde{y}),\tilde{\phi}_{t}^{\prime}(\tilde{y}))\lesssim\frac{1}{1+|\tilde{y}|}\lesssim\frac{1}{1+|y|}.$
Thus, the elementary arguments involving the mean value theorem in the proof of [DMX15, Lemma 9.4(b)]) give us
$\frac{|\tilde{\phi}(\tilde{y}_{+})-\tilde{\phi}(\tilde{y}_{-})|}{2|y_{+}-y_{-}|}\lesssim\frac{1}{1+2|y_{+}-y_{-}|}\left[\frac{1}{1+|y_{+}|}+\frac{1}{1+|y_{-}|}\right],$
which completes the proof of part (a).
For part (b), we write
$|F(t,x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|=$ $|F(t,x,\tau,\xi)-F(t^{\prime},x,\tau,\xi)+F(t^{\prime},x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|$
$\leq$ $|F(t,x,\tau,\xi)-F(t^{\prime},x,\tau,\xi)|+|F(t^{\prime},x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|$
from which we will apply the mean value theorem in the first and second variables of $F$, which will conclude the proof. For the space increment, we compute
$|\partial_{x}F|=|i\xi F|\leq\left[\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}\right]$
Where the inequality is an application of part (a). Then the mean value theorem gives us
$|F(t^{\prime},x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|\leq|x-x^{\prime}|\left[\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}\right].$ (25)
For the time increment, we use $\partial_{t}f(t,z)=e^{tz}$ to compute
$\partial_{t}F=$ $-i\tau F+\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[\partial_{t}f(t,z_{+})-\partial_{t}f(t,z_{-})]=-i\tau F+\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[e^{tz_{+}}-e^{tz_{-}}]$
$=$ $\frac{e^{-it\tau+ix\xi}}{2i|\xi|}\left[-i\tau(f(t,z_{+})-f(t,z_{-}))+e^{tz_{+}}-e^{tz_{-}}\right]$
$=$ $\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[-i\tau(f(t,z_{+})-f(t,z_{-}))+z_{+}f(t,z_{+})-1-(z_{-}f(t,z_{+})-1)]$
$=$ $\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[(z_{+}-i\tau)f(t,z_{+})-(z_{-}-i\tau)f(t,z_{-})]$
$=$ $\frac{e^{-it\tau+ix\xi}}{2i|\xi|}[(z_{+}-i\tau)f(t,z_{+})-(z_{-}-i\tau)f(t,z_{+})+(z_{-}-i\tau)f(t,z_{+})-(z_{-}-i\tau)f(t,z_{-})]$
The fact $|\xi|=|z_{+}-z_{-}|$ gives us
$|\partial_{t}F|\leq$ $\frac{|z_{+}-z_{-}||f(t,z_{+})|}{2|z_{+}-z_{-}|}+|z_{-}-i\tau|\frac{|f(t,z_{+})-f(t,z_{-})|}{2|z_{+}-z_{-}|}$
$\leq$ $\frac{1}{2}|f(t,z_{+})|+C\frac{|f(t,z_{+})-f(t,z_{-})|}{2|z_{+}-z_{-}|}$
$=$ $\frac{1}{2}\frac{|(1-e^{-\frac{at}{2}})+e^{i(\tau+|\xi|)}|}{|i(\tau+|\xi|)+\frac{a}{2}|}+C\frac{|f(t,z_{+})-f(t,z_{-})|}{2|z_{+}-z_{-}|}$
$\lesssim$ $\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}.$
For the last inequality above, the second term is again an application of part (a), and the argument for the first term is elementary. This gives us
$|F(t,x,\tau,\xi)-F(t^{\prime},x,\tau,\xi)|\lesssim|t-t^{\prime}|\left[\frac{1}{1+\frac{1}{2}|\tau+|\xi||}+\frac{1}{1+\frac{1}{2}|\tau-|\xi||}\right],$
which finishes the proof. ∎
Lemma 4.3.
  1. (a)
    For any $a,\alpha>0$,
    $\iint_{\max(|\theta|,|\zeta|)\leq 2a^{\frac{1}{\alpha}}}\left[\frac{1}{1+\frac{1}{4}(\theta-\zeta)^{2}}+\frac{1}{1+\frac{1}{4}(\theta+\zeta)^{2}}\right]|\theta-\zeta|d\theta d\zeta\lesssim Ca^{\frac{1}{\alpha}}.$
  2. (b)
    For sufficiently large $b$,
    $\iint_{\max(|\theta|,|\zeta|)>b^{\frac{1}{\alpha}},r>0}\left[\frac{1}{1+\frac{1}{4}(\theta-\zeta)^{2}}+\frac{1}{1+\frac{1}{4}(\theta+\zeta)^{2}}\right]\frac{\zeta}{1+\zeta^{2}}d\theta d\zeta\lesssim b^{-2}.$
Proof.
This lemma is [DMX15, Lemma 9.5] with $\beta=1$. ∎
Proof of Lemma 4.1.
This theorem is an analogue of [DMX15, Lemma 9.3], and the original proof can be re-written here with usage of Lemmas 4.2 and 4.3 in the appropriate places with no changes in the current setting. ∎
Lemma 4.4.
Fix $w_{0}\geq 0$ and $0<a_{1}<a_{2}$, and consider the line segment $I$ defined by (3). Then, there exists $C>0$ such that
$\operatorname{Var}(u(t,x)\mid u(t_{1},x_{1}),\dots,u(t_{n},x_{n}))\geq C\min_{1\leq i\leq n}(|t-t_{i}|+|x-x_{i}|)$ (26)
uniformly for all $n\in\mathbb{N}_{+}$ and $(t,x),(t_{1},x_{1}),\dots,(t_{n},x_{n})\in I$ with $\max_{1\leq i\leq n}t_{i}\leq t$. In particular,
$\operatorname{Var}(u(t,x)-u(s,y))\geq C(|t-s|+|x-y|)$ (27)
uniformly for all $(t,x),(s,y)\in I$.
Proof.
Recall the basic property that
$\operatorname{Var}(X\mid X_{1},\dots,X_{n})=\inf_{c_{1},\dots,c_{n}\in\mathbb{R}}\operatorname{Var}\left(X-\sum_{i=1}^{n}c_{i}X_{i}\right)$
for any centered Gaussian vector $(X,X_{1},\dots,X_{n})$. Also, note that $|t-s|+|x-y|=\sqrt{2}|t-s|$ for any pair of points $(t,x)$, $(s,y)$ on the line segment $I$. Hence, it suffices to show the existence of $C>0$ such that
$\operatorname{Var}\left(u(t,x)-\sum_{i=1}^{n}c_{i}u(t_{i},x_{i})\right)\geq C(|t-t_{n}|+|x-x_{n}|)$
uniformly for all $n\in\mathbb{N}_{+}$, $c_{1},\dots,c_{n}\in\mathbb{R}$, and $(t,x),(t_{1},x_{1}),\dots,(t_{n},x_{n})\in I$ with $t_{1}\leq\cdots\leq t_{n}\leq t$. To this end, we define
$u_{0}(t,x,t_{n},x_{n})=\int_{0}^{w_{0}}\int_{x_{n}+t_{n}-s}^{x+t-s}G(t-s,x-y)W(ds,dy),$
$u_{1}(t,x,t_{n},x_{n})=u(t,x)-u_{0}(t,x,t_{n},x_{n}).$
Then we may use independence property of the white noise $\dot{W}$ to see that
$\operatorname{Var}\left(u(t,x)-\sum_{i=1}^{n}c_{i}u(t_{i},x_{i})\right)$
$=\operatorname{Var}\left(u_{1}(t,x,t_{n},x_{n})-\sum_{i=1}^{n}c_{i}u(t_{i},x_{i})\right)+\operatorname{Var}(u_{0}(t,x,t_{n},x_{n}))\geq\operatorname{Var}(u_{0}(t,x,t_{n},x_{n})).$
By Itô-Walsh isometry,
$\operatorname{Var}(u_{0}(t,x,t_{n},x_{n}))=\frac{1}{4}\int_{0}^{w_{0}}\int_{x_{n}+t_{n}-s}^{x+t-s}e^{-a(t-s)}dyds\geq C(t+x-t_{n}-x_{n}),$
since $t_{n}\leq t$ and $x_{n}\leq x$. This completes the proof. ∎
Proof of Theorem 1.1.
Fix $w_{0}\geq 0$ and $0<a_{1}<a_{2}$. Recall (5) and consider the process $\tilde{u}$ defined by
$\tilde{u}(z)=u\left(\tfrac{w_{0}+z}{\sqrt{2}},\tfrac{-w_{0}+z}{\sqrt{2}}\right),\quad z\in[a_{1},a_{2}].$ (28)
Thanks to Lemma 4.1 and (27) above, Assumptions 2.1 and 2.3 in [LX23] are satisfied for $\tilde{u}$. Therefore, we may apply Theorem 5.2 of [LX23] to obtain the desired LIL directly. ∎
Proof of Theorem 1.2.
Again, Assumptions 2.1 and 2.3 of [LX23] are satisfied for $\tilde{u}$ defined in (28). Here, we do not have the exact form of strong local nondeterminism (SLND) in Assumption 2.2 of [LX23], but (26) in Lemma 4.4 above shows that $\tilde{u}$ still satisfies one-sided SLND in the sense that
$\operatorname{Var}(\tilde{u}(z)\mid\tilde{u}(z_{1}),\dots,\tilde{u}(z_{n}))\geq C\min_{1\leq i\leq n}(z-z_{i})$ (29)
uniformly for all $n\in\mathbb{N}_{+}$ and $z,z_{1},\dots,z_{n}\in[a_{1},a_{2}]$ with $\max_{1\leq i\leq n}z_{i}\leq z$. It is enough to use (29) in place of Assumption 2.2 in [LX23] to repeat the proof of Theorem 6.1 of [LX23] (where $Q=2$) and obtain the desired exact uniform modulus of continuity result (4) with $\sqrt{4c_{2}}\leq K_{J}\leq\sqrt{2}c_{1}$, where $c_{1}$ is the constant in (13) and $c_{2}$ is the constant in (29). Both constants depend only on $(w_{0},a_{1},a_{2})$ but not on particular sub-intervals $J$ in $I$. This completes the proof. ∎
5 Proof of Theorem 1.3
We will need the following lemma, which is the analogue of [LX22, Lemma 2.4].
Lemma 5.1.
Fix $T>0$. Then
$\mathbb{E}\left[\left(u(t,x)-u(t-\tfrac{\varepsilon_{1}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}}{\sqrt{2}})-u(t-\tfrac{\varepsilon_{2}}{\sqrt{2}},x-\tfrac{\varepsilon_{2}}{\sqrt{2}})+u(t-\tfrac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}-\varepsilon_{2}}{\sqrt{2}})\right)^{2}\right]\lesssim\varepsilon_{1}\varepsilon_{2}$
uniformly for all $t\in[0,T]$, $x\in\mathbb{R}$ and $0<\varepsilon_{1},\varepsilon_{2}<t$.
Proof.
By Itô-Walsh isometry, the left hand side of the desired inequality is
$\int_{0}^{t}\int_{\mathbb{R}}\left(\Gamma(t,x)-\Gamma(t-\tfrac{\varepsilon_{1}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}}{\sqrt{2}})-\Gamma(t-\tfrac{\varepsilon_{2}}{\sqrt{2}},x-\tfrac{\varepsilon_{2}}{\sqrt{2}})+\Gamma(t-\tfrac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}-\varepsilon_{2}}{\sqrt{2}})\right)^{2}dyds.$
For this proof, we will employ the simpler notations
$\Gamma:=\Gamma(t,x),\quad\Gamma_{1}:=\Gamma(t-\tfrac{\varepsilon_{1}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}}{\sqrt{2}}),\quad\Gamma_{2}:=\Gamma(t-\tfrac{\varepsilon_{2}}{\sqrt{2}},x-\tfrac{\varepsilon_{2}}{\sqrt{2}}),\text{ and }\Gamma_{12}:=\Gamma(t-\tfrac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}},x+\tfrac{\varepsilon_{1}-\varepsilon_{2}}{\sqrt{2}}).$
In the following, we define some subsets of the open half plane $\set{y\in\mathbb{R},s>0}$ given $x\in\mathbb{R},0<\varepsilon_{1},\varepsilon_{2}<t$ from the statement of the lemma.
[…]
We now list some useful observations about the above sets.
  1. 1.
    $R$ is a disjoint union of $R_{1},R_{2},R_{12}$, and $\tilde{R}_{12}$. It is the support of $\Gamma$ and contains the support of $\Gamma_{1},\Gamma_{2}$, and $\Gamma_{12}$.
  2. 2.
    $\tilde{R}_{12}$ has area $\varepsilon_{1}\varepsilon_{2}$ and only $\Gamma$ is nonzero on it.
  3. 3.
    $R_{1}$ has area smaller than $\varepsilon_{2}T$ and only $G$ and $\Gamma_{1}$ are nonzero on it. Similarly, $R_{2}$ has area smaller than $\varepsilon_{1}T$ and only $\Gamma$ and $\Gamma_{2}$ are nonzero on it.
  4. 4.
    $R_{12}$ has area no larger than $\frac{1}{2}(t-\varepsilon_{2}-\varepsilon_{2})^{2}<T^{2}$ and $\Gamma,\Gamma_{1},\Gamma_{2}$ and $\Gamma_{12}$ are all nonzero on it.
By the above observations, we can write the left hand side of our estimate as $\tilde{I}_{12}+I_{1}+I_{2}+I_{12}$, where
$\tilde{I}_{12}:=$ $\iint_{\tilde{R}_{12}}\Gamma^{2}dyds,\quad I_{1}:=\iint_{R_{1}}(\Gamma-\Gamma_{1})^{2}dyds,$
$I_{2}:=$ $\iint_{R_{2}}(\Gamma-\Gamma_{2})^{2}dyds,\text{ and }I_{12}:=\iint_{R_{12}}(\Gamma-\Gamma_{1}-\Gamma_{2}+\Gamma_{12})^{2}dyds.$
We will now show that all the integrals above are $\lesssim\varepsilon_{1}\varepsilon_{2}$ with the constant possibly depending on $T$, which will conclude the proof.
For $\tilde{I}_{12}$, 11-$\infty$ Hölder with the boundedness of the integrand and the area of $\tilde{R}_{12}$ gives the result.
For $I_{1}$, we first apply 11-$\infty$ Hölder, giving us
$I_{1}$ $\leq\varepsilon_{2}T\max_{s\in[0,t-\frac{\varepsilon_{1}}{\sqrt{2}}]}\left|e^{-\frac{a}{2}(t-s)}-e^{-\frac{a}{2}(t-\frac{\varepsilon_{1}}{\sqrt{2}}-s)}\right|^{2}$
$\lesssim\varepsilon_{2}\left|1-e^{-\frac{a\varepsilon_{1}}{2\sqrt{2}}}\right|\lesssim\varepsilon_{1}\varepsilon_{2}.$
A similar argument works for $I_{2}$. This leaves us with $I_{12}$, for which we have
$I_{12}$ $\leq T^{2}\max_{s\in[0,t-\frac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}}]}\left|e^{-\frac{a}{2}(t-s)}+e^{-\frac{a}{2}(t-\frac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}}-s)}-(e^{-\frac{a}{2}(t-\frac{\varepsilon_{1}}{\sqrt{2}}-s)}+e^{-\frac{a}{2}(t-\frac{\varepsilon_{2}}{\sqrt{2}}-s)})\right|$
$\lesssim\left|1+e^{\frac{a}{2}(\frac{\varepsilon_{1}+\varepsilon_{2}}{\sqrt{2}})}-(e^{\frac{a}{2}(\frac{\varepsilon_{1}}{\sqrt{2}})}+e^{\frac{a}{2}(\frac{\varepsilon_{2}}{\sqrt{2}})})\right|.$
If one Taylor expands the exponential function, one can easily see that the term inside the last absolute value is $\frac{a^{2}}{2}\varepsilon_{2}\varepsilon_{2}+O(\varepsilon_{1}\varepsilon_{2})$. This finishes the proof. ∎
Proof of Theorem 1.3.
Fix $N>0$ and $0<a_{1}<a_{2}$. It suffices to show that the existence of $0<K_{2}<\infty$ such that for all $z_{0}\in[a_{1},a_{2}]$,
$\mathbb{P}\left\{\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w+z_{0}}{\sqrt{2}}+h,\frac{-w+z_{0}}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w+z_{0}}{\sqrt{2}},\frac{-w+z_{0}}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}\leq K_{2}\text{ for all $w\in[0,N]$}\right\}=1.$
Let $0<K<\infty$ be a number whose value will be chosen later. For each integer $n\geq 2$, consider the event
$A_{n}=\left\{\sup_{w\in[0,b]}\sup_{h\in[2^{-n-1},2^{-n}]}\frac{|u\big{(}\frac{w+z_{0}+h}{\sqrt{2}},\frac{-w+z_{0}+h}{\sqrt{2}}\big{)}-u\big{(}\frac{w+z_{0}}{\sqrt{2}},\frac{-w+z_{0}}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}>K\right\}.$
For simplicity, write
$\tilde{u}(w,z)=u\big{(}\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}\big{)}.$
Define $I_{n}=[0,N]\times[2^{-n-1},2^{-n}]$ and consider the process
$X(w,h)=\tilde{u}(w,z_{0}+h)-\tilde{u}(w,z_{0}),\quad(w,h)\in I_{n}.$
Define the metric $d_{X}$ on $I_{n}$ by
$d_{X}((w,h),(w^{\prime},h^{\prime}))=\|X(w,h)-X(w^{\prime},h^{\prime})\|_{2}=(\mathbb{E}|X(w,h)-X(w^{\prime},h^{\prime})|^{2})^{1/2}.$
Let $N(I_{n},r)$ be the entropy number, i.e., the smallest number of $d_{X}$-balls of radius $r$ needed to cover $I_{n}$. Suppose $h^{\prime}<h$. Then, by triangle inequality, Proposition 2.5, and Lemma 5.1,
$d_{X}((w,h),(w^{\prime},h^{\prime}))=\|\tilde{u}(w,z_{0}+h)-\tilde{u}(w,z_{0})-\tilde{u}(w^{\prime},z_{0}+h^{\prime})+\tilde{u}(w^{\prime},z_{0})\|_{2}$
$\leq\|\tilde{u}(w,z_{0}+h)-\tilde{u}(w,z_{0}+h^{\prime})\|_{2}+\|\tilde{u}(w,z_{0}+h^{\prime})-\tilde{u}(w,z_{0})-\tilde{u}(w^{\prime},z_{0}+h^{\prime})+\tilde{u}(w^{\prime},z_{0})\|_{2}$
$\lesssim\sqrt{|h-h^{\prime}|}+\sqrt{h^{\prime}|w-w^{\prime}|}\leq\sqrt{|h-h^{\prime}|}+\sqrt{2^{-n}|w-w^{\prime}|}.$
In particular, this implies that the $d_{X}$-diameter $D_{n}$ of $I_{n}$ satisfies
$D_{n}\lesssim 2^{-n/2}$
and the entropy number satisfies
$N(I_{n},r)\lesssim\frac{N2^{-n-1}}{2^{n}r^{4}}\lesssim\left(\frac{1}{2^{n/2}r}\right)^{4}.$
Thank to Borell’s inequality [Bor75], we know that for all $v>0$,
$\mathbb{P}\left\{\sup_{(w,h)\in I_{n}}|X(w,h)|-\mathbb{E}\left[\sup_{(w,h)\in I_{n}}|X(w,h)|\right]>v\right\}\leq\exp\left(-\frac{v^{2}}{2\sup_{(w,h)\in I_{n}}\mathbb{E}[|X(w,h)|^{2}]}\right).$
In particular, by Dudley’s entropy theorem [Dud67], there exists a universal constant $C_{1}$ such that
$\mathbb{E}\left[\sup_{(w,h)\in I_{n}}|X(w,h)|\right]\leq C_{1}\int_{0}^{D_{n}}\sqrt{\log N(I_{n},r)}dr$
$\leq 4C_{1}\int_{0}^{C_{2}2^{-n/2}}\sqrt{\log(\frac{C_{2}}{2^{n/2}r})}dr=4C_{1}\int_{0}^{\infty}2C_{2}2^{-n/2}x^{2}e^{-x^{2}}dx\leq C_{3}2^{-n/2}$
for some constants $C_{2},C_{3}$ that do not depend on $n$. Moreover, by Proposition 2.5, there exists $C_{4}>0$ such that for all $n\geq 2$,
$\sup_{(w,h)\in I_{n}}\mathbb{E}[|X(w,h)|^{2}]\leq C_{4}2^{-n}.$
Therefore, if $n$ is sufficiently large so that $\frac{1}{2}K\sqrt{2^{-n-1}\log\log 2^{n}}>C_{3}2^{-n/2}$, then
$\mathbb{P}(A_{n})$ $\leq\mathbb{P}\left\{\sup_{(w,h)\in I_{n}}|X(w,h)|>K\sqrt{2^{-n-1}\log\log 2^{n}}\right\}$
$\leq\mathbb{P}\left\{\sup_{(w,h)\in I_{n}}|X(w,h)|-\mathbb{E}\left[\sup_{(w,h)\in I_{n}}|X(w,h)|\right]>\frac{1}{2}K\sqrt{2^{-n-1}\log\log 2^{n}}\right\}$
$\leq\exp\left(-\frac{\frac{1}{4}K^{2}2^{-n-1}\log\log 2^{n}}{2\sup_{(w,h)\in I_{n}}\mathbb{E}[|X(w,h)|^{2}]}\right)\leq\exp\left(-\frac{K^{2}\log\log 2^{n}}{16C_{4}}\right).$
This ensures the existence of some large constant $K<\infty$ such that $\sum_{n=2}^{\infty}\mathbb{P}(A_{n})<\infty$. This together with the Borel-Cantelli lemma implies the desired result. ∎
6 Proof of Theorem 1.4
Fix $0\leq w_{0}<w_{0}^{\prime}$, $0<z_{0}<z_{0}^{\prime}$, and let $t_{0}=w_{0}/\sqrt{2}$. We decompose $u$ as
$u(t,x)=u_{1}(t,x)+u_{2}(t,x)$
for any $(t,x)\in[t_{0},\infty)\times\mathbb{R}$, where
$u_{1}(t,x)=\int_{0}^{t_{0}}\int_{\mathbb{R}}\Gamma(t-s,x-y)W(ds,dy)=\frac{1}{2}\int_{0}^{t_{0}}\int_{x-(t-s)}^{x+(t-s)}e^{-a(t-s)/2}W(ds,dy),$
$u_{2}(t,x)=\int_{t_{0}}^{t}\int_{\mathbb{R}}\Gamma(t-s,x-y)W(ds,dy)=\frac{1}{2}\int_{t_{0}}^{t}\int_{x-(t-s)}^{x+(t-s)}e^{-a(t-s)/2}W(ds,dy).$
Note that $u_{1}$ and $u_{2}$ are independent processes and $u_{2}$ is independent of $\mathscr{F}_{t_{0}}$. Let $w\in[w_{0},w_{0}^{\prime}]$, $z\in[z_{0},z_{0}^{\prime}]$, and write $(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}})$. In order to analyze the existence and propagation of singularities, we write
$u_{1}(t+h,x+h)-u_{1}(t,x)$
$=u_{1}\left(\tfrac{w+z}{\sqrt{2}}+h,\tfrac{-w+z}{\sqrt{2}}+h\right)-u_{1}\left(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}\right)$
$=\frac{1}{2}\int_{0}^{t_{0}}\int_{x-(t-s)}^{x+(t-s)}(e^{-a(t+h-s)/2}-e^{-a(t-s)/2})W(ds,dy)+\frac{1}{2}\int_{0}^{t_{0}}\int_{x+(t-s)}^{x+(t-s)+2h}e^{-a(t+h-s)/2}W(ds,dy)$
$=\frac{e^{-at/2}(e^{-ah/2}-1)}{2}\int_{0}^{t_{0}}\int_{-\sqrt{2}w+s}^{\sqrt{2}z-s}e^{as/2}W(ds,dy)+\frac{e^{-a(t+h)/2}}{2}\int_{0}^{t_{0}}\int_{\sqrt{2}z-s}^{\sqrt{2}z-s+2h}e^{as/2}W(ds,dy)$
$=\frac{e^{-\frac{a(w+z)}{2\sqrt{2}}}(e^{-ah/2}-1)}{2}X(w,z)+\frac{e^{-\frac{a(w+z)}{2\sqrt{2}}}e^{-ah/2}}{2}(Y(z+\sqrt{2}h)-Y(z)),$ (30)
where
$X(w,z)=\int_{-\sqrt{2}w+s}^{\sqrt{2}z-s}e^{as/2}W(ds,dy),\qquad Y(z)=\int_{0}^{t_{0}}\int_{-s}^{\sqrt{2}z-s}e^{as/2}W(ds,dy)$ (31)
We first establish some basic properties for the processes $X$ and $Y$.
Lemma 6.1.
$X=\{X(w,z)\}_{(w,z)\in[w_{0},w_{0}^{\prime}]\times[z_{0},z_{0}^{\prime}]}$ is a centered Gaussian random field with
$\operatorname{Var}(X(w,z)-X(w^{\prime},z^{\prime}))\lesssim|w-w^{\prime}|+|z-z^{\prime}|$ (32)
uniformly for all $(w,z),(w^{\prime},z^{\prime})\in[w_{0},w_{0}^{\prime}]\times[z_{0},z_{0}^{\prime}]$. Hence $X$ is a.s. continuous on $w_{0},w_{0}^{\prime}]\times[z_{0},z_{0}^{\prime}]$.
Proof.
Consider $w>w^{\prime}$ in $w_{0},w_{0}^{\prime}]$ and $z,z^{\prime}$ in $z_{0},z_{0}^{\prime}]$. Assuming that $z>z^{\prime}$, we have
$\operatorname{Var}(X(w,z)-X(w^{\prime},z^{\prime}))$
$\leq 2\operatorname{Var}(X(w,z)-X(w,z^{\prime}))+2\operatorname{Var}(X(w,z^{\prime})-X(w^{\prime},z^{\prime}))$
$=2\int_{0}^{t_{0}}\int_{\sqrt{2}z^{\prime}-s}^{\sqrt{2}z-s}e^{as}dyds+2\int_{0}^{t_{0}}\int_{-\sqrt{2}w+s}^{-\sqrt{2}w^{\prime}+s}e^{as}dyds$
$=2\sqrt{2}(z-z^{\prime})\int_{0}^{t_{0}}e^{as}ds+2\sqrt{2}(w-w^{\prime})\int_{0}^{t_{0}}e^{as}ds\lesssim|z-z^{\prime}|+|w-w^{\prime}|.$
The case of $z<z^{\prime}$ is similar. This proves (32). The continuity of $X$ follows from a standard application of the Kolmogorov continuity theorem. ∎
Lemma 6.2.
$\{C_{0}^{-1}Y(z)\}_{z\geq 0}$ is a standard Brownian motion, where $C_{0}=\sqrt{2}\int_{0}^{t_{0}}e^{as}ds$.
Proof.
For any $0\leq z<z^{\prime}$, by Itô-Walsh isometry,
$\operatorname{Var}(Y(z^{\prime})-Y(z))=\int_{0}^{t_{0}}\int_{\sqrt{2}z-s}^{\sqrt{2}z^{\prime}-s}e^{as}dyds=\sqrt{2}(z^{\prime}-z)\int_{0}^{t_{0}}e^{as}ds=C_{0}(z^{\prime}-z).$
By the Kolmogorov continuity theorem, $Y$ is a.s. continuous. This shows that $B(z):=C_{0}^{-1}Y(z)$ is a centered, continuous Gaussian process with $B(0)=0$ and $\operatorname{Var}(B(z^{\prime})-B(z))=|z^{\prime}-z|$, and hence is a standard Brownian motion. ∎
We are now ready to prove Theorem 1.4. As before, in order to simplify notations, we write
$\ell(w,z)=\limsup_{h\to 0^{+}}\frac{|u\big{(}\frac{w+z}{\sqrt{2}}+h,\frac{-w+z}{\sqrt{2}}+h\big{)}-u\big{(}\frac{w+z}{\sqrt{2}},\frac{-w+z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}},$
$\ell_{i}(w,z)=\limsup_{h\to 0^{+}}\frac{|u_{i}\big{(}\frac{w+z}{\sqrt{2}}+h,\frac{-w+z}{\sqrt{2}}+h\big{)}-u_{i}\big{(}\frac{w+z}{\sqrt{2}},\frac{-w+z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}},\quad i=1,2.$
Proof of Theorem 1.4.
Fix an integer $N\in\mathbb{N}_{+}$, and take $w_{0}^{\prime}=w_{0}+N$. The processes $X$ and $Y$ defined in (31) are both $\mathscr{F}_{t_{0}}$-measurable. By Lemma 6.1,
$\mathbb{P}\left\{\sup_{w\in[w_{0},w_{0}+N]}\sup_{z\in[z_{0},z_{0}^{\prime}]}|X(w,z)|<+\infty\right\}=1.$
Since $|1-e^{-ah/2}|\lesssim h$ for $h>0$ small, the preceding implies that
$\mathbb{P}\left\{\limsup_{h\to 0^{+}}\sup_{(w,z)\in[w_{0},w_{0}+N]\times[z_{0},z_{0}^{\prime}]}\frac{e^{-\frac{a(w+z)}{2\sqrt{2}}}|e^{-ah/2}-1||X(w,z)|}{\sqrt{h\log\log(1/h)}}=0\right\}=1.$ (33)
According to Lemma 6.2, $C_{0}^{-1}Y(z)$ is a standard Brownian motion. Hence, Lévy’s modulus of continuity theorem implies that
$\limsup_{h\to 0^{+}}\sup_{z\in[z_{0},z_{0}^{\prime}]}\frac{|Y(z+\sqrt{2}h)-Y(z)|}{\sqrt{h\log(1/h)}}=C_{0}2^{3/4}\quad\text{a.s.}$ (34)
With (34), we may use Meyer’s section theorem and an argument with nested intervals as in [LX22] to obtain an $\mathscr{F}_{t_{0}}$-measurable random variable $Z$ such that $Z\in[z_{0},z_{0}^{\prime}]$ a.s. and satisfies
$\mathbb{P}\left\{\limsup_{h\to 0^{+}}\frac{|Y(Z+\sqrt{2}h)-Y(Z)|}{\sqrt{h\log\log(1/h)}}=+\infty\right\}=1.$ (35)
By (30) and (21), for all $w\in[w_{0},w_{0}+N]$,
$\begin{split}&\limsup_{h\to 0^{+}}\frac{|u_{1}\big{(}\frac{w+Z}{\sqrt{2}}+h,\frac{-w+Z}{\sqrt{2}}+h\big{)}-u_{1}\big{(}\frac{w+Z}{\sqrt{2}},\frac{-w+Z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}\\ &\geq e^{-\frac{a(w+Z)}{2\sqrt{2}}}\limsup_{h\to 0^{+}}\frac{e^{-ah/2}|Y(Z+\sqrt{2}h)-Y(Z)|}{2\sqrt{h\log\log(1/h)}}-\limsup_{h\to 0^{+}}\frac{e^{-\frac{a(w+Z)}{2\sqrt{2}}}|e^{-ah/2}-1||X(w,Z)|}{2\sqrt{h\log\log(1/h)}}.\end{split}$ (36)
In particular, the above holds for $w=w_{0}$. Hence, we may apply (33) and (35) to deduce that
$\mathbb{P}\left\{\ell_{1}(w_{0},Z)=+\infty\right\}=1.$ (37)
On the other hand, observe that
$\{u_{2}(t_{0}+t,x)\}_{(t,x)\in[0,\infty)\times\mathbb{R}}\text{ has the same law as }\{u(t,x)\}_{(t,x)\in[0,\infty)\times\mathbb{R}}.$ (38)
Indeed, since both of them are centered Gaussian random fields, this can be verified easily by comparing covariance:
$\mathbb{E}[u_{2}(t_{0}+t,x)u_{2}(t_{0}+t^{\prime},x^{\prime})]=\mathbb{E}[u(t,x)u(t^{\prime},x^{\prime})]\quad\text{for all $(t,x),(t^{\prime},x^{\prime})\in[0,\infty)\times\mathbb{R}$.}$
Hence, it follows from Theorem 1.3 that there exists $K_{2}<\infty$ such that for all fixed $z\in[z_{0},z_{0}^{\prime}]$,
$\mathbb{P}\left\{\ell_{2}(w_{0},z)\leq K_{2}\right\}=\mathbb{P}\left\{\ell(0,z)\leq K_{2}\right\}=1.$
But since $u_{2}$ is independent of $\mathscr{F}_{t_{0}}$ and the random variable $Z$ is $\mathscr{F}_{t_{0}}$-measurable, the preceding and conditioning yield
$\mathbb{P}\left\{\ell_{2}(w_{0},Z)\leq K_{2}\right\}=\mathbb{E}\left[\mathbb{P}\left\{\ell_{2}(w_{0},Z)\leq K_{2}\,\big{|}\,Z\right\}\right]=1.$ (39)
Thanks to (21), we may combine (37) and (39) to obtain
$\ell(w_{0},Z)\geq\ell_{1}(w_{0},Z)-\ell_{2}(w_{0},Z)\geq+\infty\quad\text{a.s.}$
This proves part (i) of Theorem 1.4.
To prove part (ii), let $Z$ be an $\mathscr{F}_{t_{0}}$-measurable random variable such that
$\mathbb{P}\left\{Z\in[z_{0},z_{0}^{\prime}]\text{ and }\ell(w_{0},Z)=+\infty\right\}=1.$ (40)
By Theorem 1.3 and (38), there exists $K_{2}<\infty$ such that for any fixed $z\in[z_{0},z_{0}^{\prime}]$,
$\mathbb{P}\left\{\ell_{2}(w,z)\leq K_{2}\text{ for all $w\in[w_{0},w_{0}+N]$}\right\}=\mathbb{P}\left\{\ell(w,z)\leq K_{2}\text{ for all $w\in[0,N]$}\right\}=1.$ (41)
Since $u_{2}$ is independent of $\mathscr{F}_{t_{0}}$ and $Z$ is $\mathscr{F}_{t_{0}}$-measurable, we may use conditioning and apply (41) to deduce that
$\begin{split}&\mathbb{P}\left\{\ell_{2}(w,Z)\leq K_{2}\text{ for all $w\in[w_{0},w_{0}+N]$}\right\}\\ &=\mathbb{E}\left[\mathbb{P}\left\{\ell_{2}(w,Z)\leq K_{2}\text{ for all $w\in[w_{0},w_{0}+N]$}\,\big{|}\,Z\right\}\right]=1.\end{split}$ (42)
In particular, (40) and (LABEL:E:u_2:sim:lil:Z) implies that
$\mathbb{P}\left\{\ell_{1}(w_{0},Z)=+\infty\right\}=1.$ (43)
Now, we apply (30) and (21) to obtain the reverse inequality to (LABEL:u_1:limsup:ineq): For all $w\in[w_{0},w_{0}+N]$,
$e^{-\frac{a(w+Z)}{2\sqrt{2}}}\limsup_{h\to 0^{+}}\frac{e^{-ah/2}|Y(Z+\sqrt{2}h)-Y(Z)|}{2\sqrt{h\log\log(1/h)}}$
$\geq\limsup_{h\to 0^{+}}\frac{|u_{1}\big{(}\frac{w+Z}{\sqrt{2}}+h,\frac{-w+Z}{\sqrt{2}}+h\big{)}-u_{1}\big{(}\frac{w+Z}{\sqrt{2}},\frac{-w+Z}{\sqrt{2}}\big{)}|}{\sqrt{h\log\log(1/h)}}-\limsup_{h\to 0^{+}}\frac{e^{-\frac{a(w+Z)}{2\sqrt{2}}}|e^{-ah/2}-1||X(w,Z)|}{2\sqrt{h\log\log(1/h)}}.$
Thanks to (33) and (43), we have
$\mathbb{P}\left\{e^{-\frac{a(w_{0}+Z)}{2\sqrt{2}}}\limsup_{h\to 0^{+}}\frac{e^{-ah/2}|Y(Z+\sqrt{2}h)-Y(Z)|}{2\sqrt{h\log\log(1/h)}}=+\infty\right\}=1$
and hence
$\mathbb{P}\left\{e^{-\frac{a(w+Z)}{2\sqrt{2}}}\limsup_{h\to 0^{+}}\frac{e^{-ah/2}|Y(Z+\sqrt{2}h)-Y(Z)|}{2\sqrt{h\log\log(1/h)}}=+\infty\text{ for all $w\in[w_{0},N]$}\right\}=1.$
The preceding together with (LABEL:u_1:limsup:ineq) and (33) implies that
$\mathbb{P}\left\{\ell_{1}(w,Z)=+\infty\text{ for all $w\in[w_{0},w_{0}+N]$}\right\}=1.$
Combining this with (LABEL:E:u_2:sim:lil:Z) yields
$\mathbb{P}\left\{\ell(w,Z)=+\infty\text{ for all $w\in[w_{0},w_{0}+N]$}\right\}=1.$
Since the integer $N$ can be arbitrarily large, this completes the proof. ∎

### v2

@@
 Abstract
-For the $1+1$ dimensional damped stochastic Klein-Gordon equation, we show that random singularities associated with the law of the iterated logarithm exist and propogate in the same way as the stochastic wave equation.
-This provides evidence for possible connections to microlocal analysis, ie.
-the exact regularity and singularities described in this paper should admit wavefront set type descriptions whose propagation is determined by the highest order terms of the linear operator.
+For the $1+1$ dimensional damped stochastic Klein-Gordon equation, we show that random singularities associated with the law of the iterated logarithm exist and propagate in the same way as the stochastic wave equation.
+This provides evidence for possible connections to microlocal analysis, i.e., the exact regularity and singularities described in this paper should admit wavefront set type descriptions whose propagation is determined by the highest order terms of the linear operator.
 Despite the results being exactly the same as those of the wave equation, our proofs are significantly different than the proofs for the wave equation.
@@
 Even after this simplification, many important parts of the proof are significantly different than (and we think are more intuitive from the PDE viewpoint compared to) existing proofs for the wave equation.
+2020 Mathematics Subject Classification: 60H15, 60G17.
+Keywords: stochastic partial differential equation, Klein-Gordon equation, singularity, law of the iterated logarithm
 1 Introduction
@@
 Theorem 1.4 below justifies the existence of these random singularities and shows that they propagate along the other characteristic direction.
-One of the ingredients for proving Theorem 1.4 is the following simultaneous upper bound for the law of the iterated logarithm.
-For notational convenience, we introduce a change of coordinates via a rotation by $45^{\circ}$:
+One of the ingredients for proving Theorem 1.4 is the following simultaneous upper bound for the law of the iterated logarithm. For notational convenience, we introduce a change of coordinates:
 $(w,z)=\left(\tfrac{t-x}{\sqrt{2}},\tfrac{t+x}{\sqrt{2}}\right),\qquad(t,x)=\left(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}\right).$ (5)
@@
 Remark 1.5.
+The statements of Theorem 1.1 and 1.2 as well as part (i) of Theorem 1.4 can be extended to statements about increments in all directions as opposed to only the characteristic directions.
+To not make the paper unnecessarily complicated, we chose to avoid pursuing these extensions.
+These extensions of Theorems 1.11.2 together imply that with probability 1, the $B^{s}_{\infty,\infty}$-Besov wavefront set [DRS22] of $u$ is empty if $s<\frac{1}{2}$ and is equal to $(\mathbb{R}_{+}\times\mathbb{R})\times(\mathbb{R}^{2}\setminus\{(0,0)\})$ if $s\geq\frac{1}{2}$.
+With some extra work, the same statements can be said about the more familiar $H^{s}$-Sobolev wavefront sets [Hin25, Definition 7.35, Proposition 7.36] of $u$ for the same ranges of $s$.
+We will discuss the relevance of this in the literature review.
+Remark 1.6.
 In general one can define $\mathop{}\!\mathbin{\Box}=b^{2}\partial_{tt}-c^{2}\partial_{xx}$, but a simple change of coordinates would allow one to produce the above results (with appropriately tilted characteristic lines) with the same calculations.
+In particular, what one would see is that singularities where the increments are observed in one characteristic direction would be propagated forward in time in the other characteristic direction.
+If one studies the proof of Theorem 1.3 and the propagation part of 1.4, one will discover that the proofs would fail if one tries to repeat it for any other pair of unit vectors that are not the characteristic pair.
+This is also predictable from the known behavior of hyperbolic equations, and we will discuss more of its relevance in the literature review.
 We also note that all of our results can apply to the excited stochastic Klein-Gordon equation $(a<0)$, but with extra dependence on $T$ the time horizon in all of the constants in the estimates.
-Remark 1.6.
-Theorems 1.11.2 together imply that the $B^{s}_{\infty,\infty}$-Besov wavefront set [DRS22] of $u$ is empty if $s<\frac{1}{2}$ and is equal to $(\mathbb{R}_{+}\times\mathbb{R})\times(\mathbb{R}^{2}\setminus\{(0,0)\})$ if $s\geq\frac{1}{2}$.
-With some extra work, the same statements can be said about the more familiar $H^{s}$-Sobolev wavefront sets [Hin25, Definition 7.35, Proposition 7.36] of $u$ with the same regularity parameters.
-We will discuss the relevance of this in the literature review.
 1.1 Literature review and motivation
@@
 […]
-$(\partial_{t}-\mathop{}\!\mathbin{\bigtriangleup}-m^{2})v(t,x)=\dot{W}(t,x)$ (9)
+$(\partial_{t}-\mathop{}\!\mathbin{\bigtriangleup})v(t,x)=\dot{W}(t,x)$ (9)
 will not propagate its singularities, which are known to exist in $1+1$ dimensions [HL25]. Our main results provide heuristics for this in $1+1$ dimensions using the Smolchowski- Kramers approximation
-$(\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx}-m^{2})v_{\mu}(t,x)=\dot{W}(t,x),\quad\mu\in(0,1]$ (10)
+$(\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx})v_{\mu}(t,x)=\dot{W}(t,x),\quad\mu\in(0,1]$ (10)
 where one will see (9) as $\mu\downarrow 0$ (we refer interested readers to [CF06, CS16, CX23] for the stability of this approximation and related topics).
@@
 The effect of tilting the characteristic lines as $\mu$ gets smaller is that for any singularity of $v_{\mu}$ at some point $(t_{0},x_{0})$, the propagated singularity of part (ii) of Theorem 1.4 is at $(t_{0}+\varepsilon,x_{0}+\frac{\varepsilon}{\mu})$ at time $t_{0}+\varepsilon$, which notably goes to infinity as $\mu\downarrow 0$.
-This is consistent with the PDE heuristic that hyperbolic operators have finite speed of propagation while parabolic operators have infinite speed of propagation, so $\mathop{}\!\mathbin{\Box}_{\mu}:=\mu^{2}\partial_{tt}-\partial_{xx}$ indexes a family of hyperbolic operators whose speed of propagation goes to infinity as $\mu\downarrow 0$.
+This is consistent with the PDE heuristic that hyperbolic operators have finite speed of propagation while parabolic operators have infinite speed of propagation, so $\mathcal{L}_{\mu}:=\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx}$ indexes a family of hyperbolic operators whose speed of propagation goes to infinity as $\mu\downarrow 0$.
 1.2 Paper organization and notation
@@
 $\ell_{i}(w,z)$ $=\limsup_{h\to 0^{+}}\frac{|u_{i}\big(\frac{w+z}{\sqrt{2}}+h,\frac{-w+z}{\sqrt{2}}+h\big)-u_{i}\big(\frac{w+z}{\sqrt{2}},\frac{-w+z}{\sqrt{2}}\big)|}{\sqrt{h\log\log(1/h)}},\quad i=C,L.$
-Fix $z_{0}>0$ and assume that Theorem 1.3 holds for $u_{C}$. This implies that there exists a constant $0<K_{2}<\infty$ such that $\mathbb{P}(A_{N})=1$ for all $N\in\mathbb{N}_{+}$, where
-$A_{N}=\left\{\ell_{C}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\text{ for all $w\in[0,N]$}\right\}.$
+Fix $N>0$, $0<a_{1}<a_{2}$ and assume that Theorem 1.3 holds for $u_{C}$.
+This implies that there exists a constant $0<K_{2}<\infty$ such that $\mathbb{P}(A_{z_{0}})=1$ for all $z_{0}\in[a_{1},a_{2}]$, where
+$A_{z_{0}}=\left\{\ell_{C}(w,z_{0})\leq K_{2}\text{ for all $w\in[0,N]$}\right\}.$
 An application of Lemma 3.1 with
-$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[0,N],z=z_{0}\right\}$
-shows that there is an event $B_{N}$ with $\mathbb{P}(B_{N})=1$ on which (20) holds. It is clear that
-$B_{N}\subset\left\{\ell_{L}(w,z_{0})=0\text{ for all $w\in[0,N]$}\right\}.$
-It follows from (17) and (21) that on the event $A_{N}\cap B_{N}$, which has probability 1, we have
-$\ell(w,z_{0})\leq\ell_{C}(w,z_{0})+\ell_{L}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\quad\text{for all $w\in[0,N]$.}$
-Finally, we may take $\bigcap_{N\in\mathbb{N}_{+}}(A_{N}\cap B_{N})$ as the desired event for which (6) holds.
+$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[0,N],z\in[a_{1},a_{2}]\right\}$
+shows that there is an event $B$ with $\mathbb{P}(B)=1$ on which (20) holds. It is clear that for any $z_{0}\in[a_{1},a_{2}]$,
+$B\subset\left\{\ell_{L}(w,z_{0})=0\text{ for all $w\in[0,N]$}\right\}.$
+It follows from (17) and (21) that on the event $A_{z_{0}}\cap B$, which has probability 1, we have
+$\ell(w,z_{0})\leq\ell_{C}(w,z_{0})+\ell_{L}(w,z_{0})\leq K_{2}\quad\text{for all $w\in[0,N]$.}$
 3.4 Reduction of Theorem 1.4
@@
 4 Proof of Theorems 1.1 and 1.2
+Thanks to the reduction in Section 3, we assume in the remainder of the paper that $m^{2}=a^{2}/2$.
 Denote by $\tilde{F}[f](\tau,\xi)$ the space-time fourier transform of an arbitrary $f(t,x)$ and $\tilde{\Gamma}(\tau,\xi):=\tilde{F}[\Gamma(\cdot,\circ)](\tau,\xi)$.
@@
 Note here that $\tilde{\phi}_{t}$ is a map $\mathbb{C}\supset\mathbb{R}+i\frac{a}{2}\to\mathbb{C}$, so we need not worry about $\tilde{y}=0$.
-However, it is easy to show using the elementary arguments that
+However, it is easy to show using elementary arguments that
 $\sup_{(t,\tilde{y})\in[0,T]\times\mathbb{R}+i\frac{a}{2}}\max(\tilde{\phi}_{t}(\tilde{y}),\tilde{\phi}_{t}^{\prime}(\tilde{y}))\lesssim\frac{1}{1+|\tilde{y}|}\lesssim\frac{1}{1+|y|}.$
@@
 $\frac{|\tilde{\phi}(\tilde{y}_{+})-\tilde{\phi}(\tilde{y}_{-})|}{2|y_{+}-y_{-}|}\lesssim\frac{1}{1+2|y_{+}-y_{-}|}\left[\frac{1}{1+|y_{+}|}+\frac{1}{1+|y_{-}|}\right],$
-which completes the proof of part (a).
-For part (b), we write
+which completes the proof of part (a). For part (b), we write
 $|F(t,x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|=$ $|F(t,x,\tau,\xi)-F(t^{\prime},x,\tau,\xi)+F(t^{\prime},x,\tau,\xi)-F(t^{\prime},x^{\prime},\tau,\xi)|$
@@
 Proof of Lemma 4.1.
-This theorem is an analogue of [DMX15, Lemma 9.3], and the original proof can be re-written here with usage of Lemmas 4.2 and 4.3 in the appropriate places with no changes in the current setting. ∎
+This theorem is an analogue of [DMX15, Lemma 9.3], and the original proof can be re-written here with usage of Lemmas 4.2 and 4.3 (with $\alpha=1/2$) in the appropriate places with no changes in the current setting.
+∎
 Lemma 4.4.
@@
 uniformly for all $n\in\mathbb{N}_{+}$, $c_{1},\dots,c_{n}\in\mathbb{R}$, and $(t,x),(t_{1},x_{1}),\dots,(t_{n},x_{n})\in I$ with $t_{1}\leq\cdots\leq t_{n}\leq t$. To this end, we define
-$u_{0}(t,x,t_{n},x_{n})=\int_{0}^{w_{0}}\int_{x_{n}+t_{n}-s}^{x+t-s}G(t-s,x-y)W(ds,dy),$
+$u_{0}(t,x,t_{n},x_{n})=\int_{0}^{w_{0}}\int_{x_{n}+t_{n}-s}^{x+t-s}\Gamma(t-s,x-y)W(ds,dy),$
 $u_{1}(t,x,t_{n},x_{n})=u(t,x)-u_{0}(t,x,t_{n},x_{n}).$
@@
 $\ell(w_{0},Z)\geq\ell_{1}(w_{0},Z)-\ell_{2}(w_{0},Z)\geq+\infty\quad\text{a.s.}$
-This proves part (i) of Theorem 1.4.
-To prove part (ii), let $Z$ be an $\mathscr{F}_{t_{0}}$-measurable random variable such that
+This proves part (i) of Theorem 1.4. To prove part (ii), let $Z$ be an $\mathscr{F}_{t_{0}}$-measurable random variable such that
 $\mathbb{P}\left\{Z\in[z_{0},z_{0}^{\prime}]\text{ and }\ell(w_{0},Z)=+\infty\right\}=1.$ (40)
@@
 Since the integer $N$ can be arbitrarily large, this completes the proof. ∎
+Acknowledgment. The second author is supported in part by a research startup fund of the Chinese University of Hong Kong, Shenzhen and the Shenzhen Peacock fund 2025TC0013.